Markhem Enterprises is expected to earn a net income of $75,000 this year. The company has 40,000 common shares outstanding. Each share is currently trading in the market at a multiple of 18 times. What should be the stock price of this company?
Answer #1

Income=$75000

Shares outstanding=40000

Earnings per share (EPS)= 1.875 (75000/40000)

PE Mutiple=18

stock price= EPS*PE multiple= 1.875*18= $33.75